OpenStudy (happykiddo):

Diff EQ. Solution of y''+9y'=0 y(0)=1 y'(0)=1 ?

OpenStudy (happykiddo):

Can't use euler-cauchy method. So only way left is just the euler method, but there's no initial value "h" to use euler.

OpenStudy (happykiddo):

I figured out how to this one, How bout this one? y''+9y'=0 y(0)=1 y'(0)=1.

OpenStudy (happykiddo):

Could someone explain the euler formula and how to use it, with... y''+9y'=0 y(0)=1 y'(0)=1.

OpenStudy (anonymous):

\[y(x)=\frac{10}{9}-\frac{e^{-9 x}}{9} \]

OpenStudy (happykiddo):

Okay.... Could you explain the logic behind your answer.

OpenStudy (happykiddo):

They want an exact solution. "Solve"

OpenStudy (anonymous):

\[y(x)\to c_2-\frac{1}{9} c_1 e^{-9 x} \ ? \]

OpenStudy (anonymous):

Used Mathematica v9 to solve the equation. I can delete all of my postings related to this problem if you wish.

OpenStudy (anonymous):

\[y \prime \prime+9y \prime =0\] integrate w.r.t.x \[y \prime+9y=c\] when x=0,y=1,y'=1 1+9(1)=c c=10 y'+9y=10 \[I.F.=e ^{\int\limits 9~dx}=e ^{9x}\] C.S. is \[y e ^{9x}=\int\limits 10~e ^{9x}dx+C\] \[ye ^{9x}=\frac{ 10 e ^{9x} }{ 9 }+C\] again~when~x=0,y=1 \[1(1)=\frac{ 10(1) }{ 9 }+C,C=1-\frac{ 10 }{ 9 }=-\frac{ 1 }{ 9 }\] C.S.is \[y e ^{9x}=\frac{ 10e ^{9x} }{ 9 }-\frac{ 1 }{ 9 }\]